# IntGrad MT: Eliciting LLMs’ Machine Translation Capabilities with Sentence Interpolation and Gradual MT

### v2

@@ Abstract
Recent Large Language Models (LLMs) have demonstrated strong performance in translation without needing to be finetuned on additional parallel corpora. However, they still underperform for low-resource language pairs. Previous works have focused on mitigating this issue by leveraging relevant few-shot examples or external resources such as dictionaries or grammar books, making models heavily reliant on these nonparametric sources of information. In this paper, we propose a novel method named IntGrad MT that focuses on fully exploiting an LLM’s inherent translation capability. IntGrad MT achieves this by constructing a chain of few-shot examples, each consisting of a source sentence and the model’s own translation, that rise incrementally in difficulty. IntGrad MT employs two techniques: Sentence Interpolation, which generates a sequence of sentences that gradually change from an easy sentence to translate to a difficult one, and Gradual MT, which sequentially translates this chain using translations of earlier sentences as few-shot examples for the translation of subsequent ones. With this approach, we observe a substantial enhancement in the xCOMET scores of various LLMs for multiple languages, especially in low-resource languages such as Hindi(8.26), Swahili(7.10), Bengali(6.97) and Marathi(13.03). Our approach presents a practical way of enhancing LLMs’ performance without extra training.
@@ 1 Introduction
Recent Large Language Models (LLMs) have shown strong performance in translation tasks without the need for fine-tuning on specific parallel datasets. Previous studies have demonstrated that LLMs’ translation capabilities are reliable in most use cases, particularly when the source and target language are high-resource languages (Zhu et al., 2024; Robinson et al., 2023; Jiao et al., 2023). However, because LLMs require training on large corpora, they still face challenges when translating low-resource languages that are not sufficiently represented in the training corpora.(Stap & Araabi, 2023; Robinson et al., 2023; Enis & Hopkins, 2024).
@@
In this paper, we examine whether we can improve LLMs’ translation capabilities without relying on external knowledge. We aim to answer this question by considering a simple fact: the machine translation task can be defined as a mapping between two (sub)spaces for the source and target language. In the source language space, there are regions where the model performs well in translation and regions where it does not. The key idea is that if we can gradually expand the areas where the model performs well by feeding it neighboring examples, we can enhance its translation capabilities in the areas where it performs poorly.
We propose IntGrad MT to achieve this by connecting sentences from the regions where the LLM performs well in translation and regions where it does not. IntGrad MT consists of two key techniques: Sentence Interpolation and Gradual MT. Sentence Interpolation is a prompting technique that generates a sequence of sentences gradually transitioning from one to another. Gradual MT is a technique in which the model iteratively translates a list of sentences, using its previous translations as few-shot examples for subsequent sentences. Through sentence interpolation, we first establish a pathway to the sentences in regions where the model performs poorly, then gradually expand the area where the model can excel, utilizing its in-context learning capabilities. The key concepts are illustrated in Figure 1.
Figure 1: Figurative illustration of IntGrad MT. Machine translation task can be defined as a mapping between two (sub)spaces for the source and target language. In the source language space, there are regions where the model performs well in translation and regions where it does not. IntGrad MT expand the areas where the model performs well by feeding LLM with neighboring examples, eventually reaching the areas where it performs poorly.
We test the effectiveness of IntGrad MT by applying it to four different LLMs — GPT-3.5 Turbo, Mistral Nemo Instruct, Llama 3.1 70B Instruct, and Llama 3.1 8B Instruct — across seven target languages: German (De), Chinese (Zh), Hindi (Hi), Korean (Ko), Swahili (Sw), Marathi (Mr), and Bengali (Bn). Our results show consistently large performance gains, particularly in low-resource languages.
Additionally, we conduct an ablation study on English-to-Korean translation using ChatGPT to determine the optimal settings in relation to the following three questions:
  • How should we select the start sentences for interpolation?
  • How should we aggregate multiple translation results from different interpolation paths?
  • To what sentences IntGradMT can be effective?
Our contributions are summarized as follows:
  • We propose IntGrad MT, a novel method for machine translation that leverages two key methods: Sentence Interpolation and Gradual MT.
  • We test the effectiveness of IntGrad MT with various LLMs and across different target languages.
  • We conduct an ablation study to identify the optimal configuration.
  • We analyze the geometrical characteristics of sentence interpolation.
@@ 2.1 Enhancing LLMs’ Translation Capabilities without Fine-tuning
Modern LLMs show high capabilities in translation tasks in high- resource languages, but not in low-resource languages (Jiao et al., 2023; Stap & Araabi, 2023; Zhu et al., 2024; Enis & Hopkins, 2024). There are several works that have focused on enhancing LLMs’ translation capabilities without additional fine-tuning. A primary strategy involves leveraging LLMs’ ability to learn from demonstrations or descriptions (Brown et al., 2020; Wei et al., 2022). Studies have explored selecting appropriate exemplars for few-shot learning and demonstrating linguistic knowledge (Agrawal et al., 2022; Vilar et al., 2023; Zhang et al., 2024), or augmenting LLMs with chains of multilingual dictionaries (Lu et al., 2024). Besides providing a demonstration or description, choosing the right temperature or prompting strategy has also been examined (Peng et al., 2023). Similar to previous research, our method aims to improve LLMs’ MT capabilities without fine-tuning, focusing instead on eliciting the models’ inherent capabilities with sentence interpolation and gradual MT rather than providing them with few-shot examples or external knowledge.
@@ 2.2 Self-Demonstration
Manually generating appropriate exemplars for in-context learning can be resource-intensive. To address this challenge, previous studies have explored enabling models to generate their own few-shot examples for tasks such as classification(Lyu et al., 2023; Kim et al., 2022) or other reasoning tasks(Zhang et al., 2023; Li et al., 2024). Our work is aligned with these efforts, as it also focuses on generating the model’s own few-shot examples. However, none of these approaches have tried to create examples such as ours, since our approach aims to produce the tailored example by gradually expanding the example set with similar yet distinct examples.
3 Methods
IntGrad MT consists of two main steps: Sentence Interpolation (§3.1) and Gradual MT (§3.2). We first introduce these two steps and present how we combine them (§3.3).
3.1 Sentence Interpolation
We propose sentence interpolation, which is a prompting technique asks model to create a list of sentences that gradually change from start sentence to end sentence.
For example, following is the prompt that we use in our experiments:
{shadedquotation}
@@ 3.1 Sentence Interpolation
The objective of this technique is to generate a list of sentences where each sentence is distinct, yet not excessively different from its adjacent sentences. By prompting the LLM to autonomously create the list, rather than mechanically altering sentences, we can obtain natural sentences that are suitable for use as in-context learning examples. We call this list of interpolated sentences the interpolation path.
In practice, we utilized three few-shot examples from GPT-4(OpenAI et al., 2024) to control the output format. See Appendices A and B for these examples and sample interpolation paths.
@@ 3.2 Gradual MT
Gradual MT is a prompting technique that lets an LLM utilize its previous translations as a prompt. This approach sequentially processes a list of sentences, translating each one while using the previous translation results as few-shot examples for the current sentence. An illustration of the Gradual MT process is shown in Figure 2. Gradual MT can effectively guide LLMs in accurately translating unfamiliar sentences when combined with sentence interpolation.
Figure 2: Illustration of Gradual MT. Gradual MT iteratively processes a list of sentences, translating each one while using the previous translation results as few-shot examples for the current sentence.
@@ 3.3 Overall Method
IntGrad MT combines sentence interpolation and gradual MT. The illustration of the algorithm can be seen in Figure 3.
Figure 3: Illustration of the IntGrad MT algorithm. IntGrad MT combines Sentence Interpolation with Gradual MT. In step 1, $k$ start sentences are selected from a pre-created start sentence pool. In step 2, start sentences are interpolated to the end sentence, creating $k$ individual interpolation paths. Each path is then processed through Gradual MT, generating translation results for each sentence. In step 3, MT results from each interpolation path are aggregated to produce a single output MT. In step 4, the sentences to which IntGrad MT is applied are selected.
Step 0: Start Sentence Pool Creation.
Before applying IntGrad MT, the start sentence pool must be created before. This pool consists of sentences that the LLM can already translate accurately. It is crucial to ensure that the source sentences included in this pool have been tested and verified, confirming the LLM can generate high-quality translations in a zero-shot setting. We refer to the source sentences in the start sentence pool as start sentences.
@@
IntGrad MT begins with selecting $k$ start sentences from the start sentence pool by calculating the similarity with the source sentence that the LLM is trying to translate, which we call an end sentence. Measuring similarity between sentences can be done in various ways. In this paper, we utilized SBERT similarity(Reimers & Gurevych, 2019) as a primary metric. Details about start sentence selection strategy can be found in Section 6.1.1.
Step 2: Sentence Interpolation & Gradual MT.
After selecting $k$ start sentences, sentence interpolation is performed between each start sentence and the corresponding end sentence, creating $k$ individual interpolation paths. The paths are then processed through the Gradual MT, generating translation results for each sentence.
Step 3: MT Results Aggregation.
After generating a list of translation, we proceed to aggregate the final translation from each path. If a single start sentence was selected in Step 1, this step is skipped. There are several methods for aggregating translations; in our approach, we input all the translation results into the LLM once again as few-shot examples to generate the final translation. A detailed explanation of this aggregation strategy can be found in Section 6.1.2. After aggregation, we obtain the final translation, which we refer to as the output.
Step 4: Output Selection.
The final step is to select the sentences to which we apply IntGrad MT. There are two motivations for not applying the method to all sentences. First, if we decide which sentences will undergo all the steps of IntGrad MT before executing them, we can substantially reduce the computing cost. Second, even if we don’t select sentences beforehand, it is still meaningful to selectively use the outcome only when it is better than the zero-shot translation result by comparing reference-free QE model scores. We tested three different strategies in our ablation study (§6.1.3): ‘Thr.’, ‘Delta’, and ‘Thr.+Delta’. ‘Thr.’ applies zero-shot translation first and uses IntGrad MT only if the QE score is below a certain threshold. ‘Delta’ chooses outputs only if they outperform zero-shot translations. ‘Thr. +Delta’ combines these two strategies.
@@ 4.1 Setup
For translation we use four different LLMs: ChatGPT(GPT-3.5-Turbo-0125) [^1] https://platform.openai.com/docs/models/gpt-3-5-turbo, Mistral Nemo(Mistral-Nemo-Instruct-2407) [^2] https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407/ and two different sized Llama 3.1 Instruct models(Llama-3.1-70B-Instruct, Llama-3.1-8B-Instruct)(Dubey et al., 2024). ChatGPT is accessed via OpenAI’s API, and the others are run locally. For sentence interpolation, we employ Qwen2-72b-Instruct (Yang et al., 2024) with quantization. See Appendices C and D for settings for translation and interpolation. For output selection, we utilize a reference-free QE model CometKiwi (Rei et al., 2023) to avoid peeking at the gold translations. CometKiwi produces a Direct Assessment (DA) score, which rates translation quality on a scale from 0 to 100, normalized to a range of 0 to 1. Lastly, we used all-mpnet-base-v2333https://huggingface.co/sentence-transformers/all-mpnet-base-v2 for SBERT sismilarity calculation.
Target Languages.
@@
We fixed English as the source language. The target languages tested in the experiments are German (De), Chinese (Zh), Korean (Ko), Hindi (Hi), Swahili (Sw), Bengali (Bn), and Marathi (Mr). Based on Joshi et al. (2020)’s 6 scale taxonomy of language resource level, we classify German and Chinese as high resource languages, Korean and Hindi as mid resource, and the rest as low resource. For the Llama 3.1 models, we conduct tests only on German and Hindi, as those models do not support other languages. Mistral Nemo Instruct does not officially support Ko, Hi, Sw, Bn, and Mr, but we conduct experiments on those languages nevertheless, as it has some capability to generate them.
Dataset.
We use the FLORES-200 benchmark dataset(Team et al., 2022) for validation and evaluation. We first utilize the dev split of the dataset to create start sentence pool. During evaluation, to test the effect of the output selection strategy, we selected 10% of the test portion of the dataset to set the DA score threshold. The remaining 90% is used to assess the overall performance on the dataset.
Start Sentence Pool Creation.
@@
The initial sentence pool is created by translating source sentences from the dev split of the FLORES-200 dataset using a zero-shot approach. Each source sentence is translated five times and evaluated with xCOMET (Guerreiro et al., 2024). xCOMET predicts DA score normalized to a range of 0 to 1. The most frequently occurring translation is selected as the representative translation. If no translation is repeated, the one with a score closest to the average is chosen. After selecting each representative translation, the top 100 translation pairs with the highest DA scores are selected.
Baselines.
We compare the outcomes with zero-shot MT and three-shot MT results. To assess the pure impact of our method, we use the same start sentences for interpolation as in the few-shot examples.
Evaluation.
We use the DA score evaluated by xCOMET as the primary metric for our evaluation, multiplying it by 100 for readability. Additionally, we employ MetricX (Juraska et al., 2023) to obtain an MQM score, which assesses translation errors on a scale from 0 to 25, with lower scores indicating better quality. Lastly, we also check the DA score evaluated by CometKiwi.
5 Results
We first conducted various ablation settings for English-Korean translation using ChatGPT to identify the optimal configuration such as start selection(step 1) and result aggregation(step 4). Subsequently, this configuration was applied to other IntGrad MT experiments utilizing different LLMs and languages. A detailed explanation of the ablation study is provided in the next section. In this section, we report xCOMET results of IntGrad MT on various LLMs and target languages in Table 1. Similar results based on MetricX and CometKiwi are presented in Appendix E.
As shown in Table 1, IntGrad MT outperforms the baselines in all models, whether large or small, and languages of diverse resource levels, with no degeneration that often occurs when reusing the model’s own output. This indicates that LLMs possess hidden multilingual and linguistic knowledge that can be exploited for machine translation with the correct techniques. Compared to the zero-shot baseline, IntGrad MT increases the xCOMET score by a maximum of 3.03, 8.26, and 13.03 points in high, mid, and low resource languages, respectively. It incurs the largest improvements for low- and mid- resource languages, suggesting that it holds the potential to improve the quality of translation for these languages that LLMs are currently unable to translate well. IntGrad MT also outperforms the 3 shot baselines by wide margins in low resource languages. Mistral Nemo’s xCOMET scores improve by 5.15, 4.61, 4.42 for Swahili, Bengali, Marathi, respectively. This supports the advantage of our method in formulating tailored prompts using the example translation pairs over simply listing them.
Table 1: xCOMET scores of IntGrad MT with the best ablation setting for different LLMs and target languages. The strategy used includes selecting the starting sentences by sorting with SBERT similarity and tree edit distance, aggregating MT results by feeding all gradual MT results into the LLM, and selecting the output MT only if its CometKiwi score is higher than the zero-shot translation. The results of setting the output selection strategy as Thr.+Delta, which selects the output MT only if the CometKiwi score of the zero-shot MT is below a certain threshold and the output MT’s score is higher than the zero-shot score, are presented below the main results as supplementary information. Zero-shot and three-shot are the baselines, with the three-shot examples selected using the same starting sentence selection method as IntGrad MT. Scores are multiplied by 100 for readability, and the first two highest scores for each MT model and language are highlighted in bold and underlined. Llama translation results for languages other than De and Hi are not available, as Llama does not support those languages. The numbers following the up arrows indicate the delta with respect to the zero-shot baseline.

### v3

# BridG MT: Enhancing LLMs’ Machine Translation Capabilities with Sentence Bridging and Gradual MT

@@ Abstract
Recent Large Language Models (LLMs) have demonstrated impressive translation performance without requiring fine-tuning on additional parallel corpora. However, they still face significant challenges in certain scenarios, particularly when translating low-resource languages. A common approach to address this issue is to provide external knowledge, such as few-shot examples, to assist LLMs in translating specific source sentences. However, this method is fundamentally limited by the quality or quantity of relevant sources, which cannot always be guaranteed. To reduce LLMs’ reliance on external sources, we propose BridG MT, a method that combines Sentence Bridging, which generates a sequence of sentences as a bridge that gradually transition from easy-to-translate to more difficult, and Gradual MT, which sequentially translates these sentences using earlier translations as few-shot examples for subsequent ones. Experiments conducted on four LLMs across seven languages demonstrate that our method effectively enhances translation performance, even outperforming translation methods that rely on a large number of few-shot examples.
[…]
[ backgroundcolor=gray!10, linecolor=gray!10, leftmargin=10pt, rightmargin=10pt, skipabove=5pt, skipbelow=5pt ]shadedquotation
BridG MT: Enhancing LLMs’ Machine Translation Capabilities
with Sentence Bridging and Gradual MT
Seungwoo Choi, Gahyun Yoo, Jay-Yoon Lee* Seoul National University {rhdn520, padme0421, lee.jayyoon}@snu.ac.kr
@@ 1 Introduction
Figure 1: Figurative illustration of BridG MT. Machine translation maps between source and target language spaces. BridG MT expands well-performing areas (blue) by leveraging neighboring examples, reaching poorly performing areas (red). Circle sizes indicate output noisiness.
Recent Large Language Models (LLMs) have shown strong performance in translation tasks without the need for fine-tuning on specific parallel datasets. Previous studies have demonstrated that LLMs’ translation capabilities are reliable in most use cases, particularly when the source and target language are high-resource languages (Zhu et al., 2024; Robinson et al., 2023; Jiao et al., 2023). However, because LLMs require training on large corpora, they still face challenges when translating low-resource languages that are not sufficiently represented in the training corpora.(Stap and Araabi, 2023; Robinson et al., 2023; Enis and Hopkins, 2024).
@@
In this paper, we investigate whether LLMs’ translation capabilities can be improved without relying on external knowledge. Our approach is grounded in a fundamental idea: machine translation can be understood as a mapping between two (sub)spaces representing the source and target languages. Within the source language space, some regions yield accurate translations, while others do not. The key question is whether we can expand the model’s confident region by refeeding examples from it, ultimately extending into weaker regions and improving performance in those areas.
We propose BridG MT to enhance machine translation by bridging regions where the LLM performs well with those where it struggles. BridG MT integrates two key techniques: Sentence Bridging and Gradual MT. Sentence Bridging is a prompting method that generates a sequence of sentences, progressively transitioning between them. Gradual MT iteratively translates a list of sentences, using the model’s previous translations as few-shot examples for subsequent ones. Sentence Bridging helps bridge high- and low-performance regions, while Gradual MT gradually expands the model’s strong performance areas by leveraging in-context learning. These concepts are visually illustrated in Figure 1.
We evaluate the effectiveness of BridG MT by applying it to four different LLMs: GPT-3.5, Mistral-Nemo-Instruct, Llama-3.1-70B-Instruct, and Llama-3.1-8B-Instruct. The experiments cover seven target languages: German (De), Chinese (Zh), Hindi (Hi), Korean (Ko), Swahili (Sw), Marathi (Mr), and Bengali (Bn). The results demonstrate that BridG MT significantly enhances the translation capabilities of LLMs, particularly in low-resource languages. To the best of our knowledge, we are the first to propose the concept of Sentence Bridging.
@@ 2.1 Enhancing LLMs’ Translation Capabilities without Fine-tuning
Modern LLMs demonstrate strong translation capabilities in high-resource languages but struggle with low-resource languages (Jiao et al., 2023; Stap and Araabi, 2023; Zhu et al., 2024; Enis and Hopkins, 2024). Several studies have focused on enhancing LLMs’ translation performance without additional fine-tuning. A primary approach involves leveraging LLMs’ ability to learn from demonstrations or descriptions through in-context learning (Brown et al., 2020; Wei et al., 2022). Researchers have explored methods such as selecting appropriate exemplars for few-shot learning and demonstrating linguistic knowledge (Agrawal et al., 2022; Vilar et al., 2023; Zhang et al., 2024). Beyond simply providing examples, some approaches conduct a prior analysis of the sentence to be translated, using LLMs to extract relevant information. These approaches then supplement the translation process with resources such as chains of multilingual dictionaries (Lu et al., 2024), or adopt human-like strategies by providing the LLM with keywords, topics, and generated demonstrations (He et al., 2024). Other studies have instead focused on generating multiple translation candidates and selecting the optimal output through ranking mechanisms such Minimum Bayes Risk decoding (Fernandes et al., 2022).
@@ 2.2 Self-Demonstration
Manually generating appropriate exemplars for in-context learning can be resource-intensive. To address this challenge, previous studies have explored enabling models to generate their own few-shot examples for tasks such as classification (Lyu et al., 2023; Kim et al., 2022) or other reasoning tasks (Zhang et al., 2023; Li et al., 2024). Our work is aligned with these efforts, as it also focuses on generating the model’s own few-shot examples. However, these techniques have yet to be applied to translation, and BridG MT’s novel approach of intentionally bridging confident regions to current predictions remains unexplored.
3 Methodology
We introduce two major components of BridG MT—Sentence Bridging and Gradual MT—in Sections 3.1 and 3.2, respectively, and then describe how we combine them in Section 3.3.
Figure 2: Illustration of the BridG MT algorithm. BridG MT integrates Sentence Bridging with Gradual MT. In Step 1, $k$ start sentences are selected from a predefined start-sentence pool. In Step 2, these start sentences are bridged to the end sentence, creating $k$ individual bridges. Each bridge is then processed through Gradual MT, generating translation results for every sentence along the path. In Step 3, the MT results from all bridges are aggregated into a single output translation. Optional pre- and post-filtering steps can be applied between Steps 1 and 2, and again after aggregation, to refine the sentences on which BridG MT is applied.
3.1 Sentence Bridging
Sentence bridging is a prompting technique that asks the model to generate a list of sentences that gradually transition from a start sentence to an end sentence. The objective of this technique is to generate a list of sentences where each sentence is distinct, yet not excessively different from its adjacent sentences. We call these sentences the bridge. In our experiments, we utilized three bridging examples from GPT-4(OpenAI et al., 2024) as a few-shot to control the output format. See Appendices A and B for these examples and sample bridges. Following is the prompt that we use in our experiments:
@@ 3.2 Gradual MT
Gradual MT is a prompting technique that enables an LLM to leverage its previous translations as prompts. This approach sequentially processes a bridge, translating each source sentence while using the preceding translation results as few-shot examples for the current sentence. An illustration of the gradual MT algorithm is provided in Algorithm 1. By applying gradual MT, we construct an expanding set of few-shot examples, which become increasingly useful for translating the final sentence that the model must process.
Algorithm 1 Pseudo Code for Gradual MT
  Input: Bridge $X=\{x_{1},x_{2},...,x_{n}\}$, $n=|X|$, Translation model $M$
  Set $fewshot=[]$
  for $x_{i}\in X$ do
     $\hat{y}_{i}=M(x_{i},fewshot)$
     Append $\{x_{i},\hat{y}_{i}\}$ to $fewshot$
  end for
  Output: $\hat{y}_{n}$
@@ 3.3 Overall Method
BridG MT combines sentence bridging and gradual MT. An illustration of the algorithm is shown in Figure 2.
Step 0: Start Sentence Pool Creation
Before applying BridG MT, the start sentence pool must first be created. This pool consists of sentences that the zero-shot LLM can translate with high accuracy. Since we are considering a scenario where there is no reliable gold translation, we utilize a reference-free quality estimation(QE) model such as CometKiwi(Rei et al., 2023) to construct the start sentence pool. We refer to the source sentences in the start sentence pool as start sentences.
@@
BridG MT begins with selecting $k$ start sentences from the start sentence pool by calculating the similarity with the source sentence that the LLM is trying to translate, which we call an end sentence. Measuring similarity between sentences can be done in various ways. In this paper, we utilized SBERT similarity(Reimers and Gurevych, 2019) as a primary metric. Details about start sentence selection strategy can be found in Section 6.1.
Step 2: Sentence Bridging & Gradual MT
After selecting start sentences, sentence bridging is performed between each start sentence and the corresponding end sentence, creating $k$ separate bridges. The bridges are then processed through the Gradual MT, generating translation results for each sentence.
Step 3: MT Results Aggregation
When the number of start sentences is more than one, we proceed to aggregate the translation results from each bridge. This step is skipped if we choose to work with single start sentence in Step 1. There could be several methods for aggregating translations; in our approach, we input all the translation results into the LLM once again as few-shot examples to generate the final translation. A detailed ablation of this aggregation strategy can be found in Section 6.2. After aggregation, we obtain the final translation, which we refer to as the output.
Pre- & Post-filtering
BridG MT can be applied to any sentence; however, it is often more effective to use it selectively for two reasons. First, determining which sentences will undergo BridG MT before its execution can significantly reduce computational costs. Second, even after applying BridG MT, it is better to discard the BridG MT result if its quality is worse than the zero-shot translation. We explored these two possibilities—referred to as Pre-filtering and Post-filtering—in Section 6.3 and show that these methos can improve quality with smaller computation than full application of BridG MT.
@@ 4.1 Setup
For translation, we use four different LLMs: GPT-3.5 (GPT-3.5-Turbo-0125) [^1] https://platform.openai.com/docs/models/gpt-3-5-turbo, Mistral-Nemo (Mistral-Nemo-Instruct-2407) [^2] https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407/, and two Llama models (Llama-3.1-70B-Instruct, Llama-3.1-8B-Instruct) (Dubey et al., 2024). ChatGPT is accessed via OpenAI’s API, while the other models ran on local GPU. For sentence bridging, we employ Qwen2-72B-Instruct (Yang et al., 2024) with 4-bit quantization as our primary bridging model. We also tested Llama-3.1-8B-Instruct and Llama-3.1-3B-Instruct to test the generalizability of the method in more cost-efficient settings. See Appendices D and E for settings for translation and bridging. For pre- and post-filtering, we utilize a reference-free QE model CometKiwi (Rei et al., 2023) to avoid peeking at the gold translations. CometKiwi predicts a DA score, which rates translation quality on a scale from 0 to 100, normalized to a range of 0 to 1. Lastly, we used all-mpnet-base-v2333https://huggingface.co/sentence-transformers/all-mpnet-base-v2 for SBERT sismilarity calculation in retrieving start sentences and few-shot examples for baseline experiments. All experiments were conducted once.
Target Languages
@@
We fixed English as the source language. The target languages tested in the experiments are German (De), Chinese (Zh), Korean (Ko), Hindi (Hi), Swahili (Sw), Bengali (Bn), and Marathi (Mr). Based on Joshi et al. (2020)’s 6 scale resource level, we classify German and Chinese as high resource, Korean and Hindi as mid resource, and the rest as low resource. For GPT 3.5, we experiment on every languages. For the Llama 3.1 models, we only experiment on German and Hindi, as they do not support other languages. Mistral-Nemo does not officially support Ko, Hi, Sw, Bn, and Mr, but we conduct experiments on those languages nevertheless, as it has some capability to generate them.
Dataset
We use the FLORES-200 benchmark dataset (Team et al., 2022) for validation and evaluation. The development split is used to construct the start sentence pool. For the test set, we sample 90% of the data and reserve the remaining 10% to determine the QE score threshold for pre-filtering. We also use NTREX-128 (Federmann et al., 2022) to evaluate whether BridG MT remains effective when the distribution of the start sentence pool differs from that of the test set. The results for NTREX-128 are presented in Section 7.
Start Sentence Pool Creation
@@
The start sentence pool is created by translating source sentences from the dev split of the FLORES-200 dataset using a zero-shot approach. Each source sentence is translated five times and evaluated with CometKiwi-a reference-free model. The most frequently occurring translation is selected as the representative translation. If no translation is repeated, the one with a score closest to the average is chosen. After selecting each representative translation, the top 100 translation pairs with the highest DA scores are selected.
Table 1: xCOMET scores of BridG MT across different translation models and target languages. Sentence bridging was performed using Qwen2-72B-Instruct. Scores are multiplied by 100 for readability. For each MT model and language, the highest score is shown in bold and the second-highest is underlined. Cells are left blank when a specific model or method does not support the corresponding language.